OpenStudy (anonymous):

If tan x° = 11 divided by r and cos x° = r divided by s, what is the value of sin x°? sin x° = s divided by 11 sin x° = 11 divided by s sin x° = 11r sin x° = 11s

OpenStudy (raden):

hint : sin x = tan x * cos x

OpenStudy (anonymous):

\[\tan x=\frac{ \sin x }{\cos x }=\frac{ 11 }{r },\sin x=\frac{ 11 }{r }*\cos x,\cos x=\frac{ r }{s }\] find sin x .

OpenStudy (anonymous):

Are r and s equal?

OpenStudy (anonymous):

\[\sin x=\frac{ 11 }{ r }*\frac{ r }{ s }=?\]